OpenStudy (anonymous):

ok i do not understand how to divide this exponent thingy -2w^3 over 5w^6v^5 times 11w^5 over 4w^3

OpenStudy (anonymous):

\[\frac{-2w^3}{5w^6v^5} \times \frac{11w^5}{4w^3}\]\[=\frac{-22w^{3+5}}{20w^{6+3}v^5}\]\[=\frac{-11w^{8}}{10w^{9}v^5}\]\[=\frac{-11w^{8-9}}{10v^5}\]\[=\frac{-11w^{-1}}{10v^5}\]\[=\frac{-11}{10v^5w}\]

OpenStudy (anonymous):